OpenStudy (anonymous):

let V be the vector space of polynomial over R with inner product defined by.....

OpenStudy (anonymous):

\[<f,g>=\int\limits_{0}^{1}f(t)g(t)dt\] Let D be the derivative operator on V, D(f)=df/dt. Show that there is no operator D* on V such that <D(f),g> = <f,D*(g)>